You decide to save up for a big trip and deposit 882 in a special vacation account today. If your account earns 4.94% annually, how much will you have available to spend in 8 years? Round to two decimal places (Ex. $0,000.00)

If you want to have $2,971 in 8 years, how much money must you put in a savings account today? Assume that the savings account pays 5.1% and it is compounded annually. Round to two decimal places (Ex. $000.00)

Your company has received a $50,000 loan from an industrial finance company. The annual payments are $6,202.70. If the company is paying 9 percent interest per year, how many loan payments must the company make?

15

13

12

19
Answer #1
a) Future value in 8 years = P*(1+i)^n Where,
= 882*(1+0.0494)^8 P $         882.00
= $ 1,297.17 i 4.94%
n 8
b) Present value = A*(1+i)^-n Where,
= 2971*(1+0.051)^-8 A $     2,971.00
= $ 1,995.63 i 5.10%
n 8
c) Number of period = =NPER(rate,pmt,pv) Where,
= 15 rate 9%
pmt $     6,202.70
pv $       -50,000